# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ Appendix B Exact solvability of QP systems from local constraint
Now we show the exact solvability for the spinful quasiperiodic system when the hopping coupling matrices satisfy the condition
$\det|\Pi_{j}|=0,\quad p_{j}^{s}=\mathrm{Const.},$ (69)
@@
which describe the effective motion with nearest neighbor hopping and the effective on-site potential is energy dependently dressed by the local modulation from the accompanied internal degree of freedom.
@@
Appendix C 1D QP mosaic lattice model
In this section, we introduce the one-dimensional quasiperiodic mosaic lattice model Wang et al. (2020a); Zhou et al. (2023) and rewrite it into our 1D spinful QP framework. The generic mosaic lattice refers that the modulation of on-site potential and hopping coefficients are in a mosaic manner, namely the the QP modulation or the uniform modulation only appears every $\kappa$ sites, as we will see in the following.
The original type-I and type-II mosaic lattice model is distinguished by the presence of alternating mosaic pattern in hopping coefficients. If the QP modulation only exhibits in the on-site potential and the hopping coefficients are always uniform or always quasiperiodic, then the mosaic models belongs to type-I mosaic model. The Hamiltonian of the type-I mosaic lattice model is given by Wang et al. (2020a):
@@
with the on-site potential given by
$V_{j}=\begin{cases}V\cos(2\pi\alpha j+\theta),&j=0\quad{\rm mod}\leavevmode\nobreak\ \kappa,\\ 0,&j\neq 0\quad{\rm mod}\leavevmode\nobreak\ \kappa,\end{cases}$ (80)
where $c_{j}$ is the annihilation operator at site $j$, and $\lambda$, $V$, and $\theta$ denote the nearest-neighbor hopping coefficient, on-site potential amplitude, and phase offset, respectively. $\alpha$ is an irrational number, and $\kappa$ is an integer. The quasiperiodic potential periodically occurs every $\kappa$ sites.
If the system exhibits mosaic modulation in hopping coefficients, then it belongs to type-II mosaic model Zhou et al. (2023) and the Hamiltonian is given by
@@
If we focus on the case $\kappa=2$, then both two moasic lattice models exhibit a even-odd sublattice structure, which can be recasted into the spinful QP chain model. We can relabel the even and odd site as spin-up and spin-down degrees of freedom and further relabel the site index, then the Hamiltonian can be rewritten as
@@
where $c_{j}=(c_{j,\uparrow},c_{j,\downarrow})^{\intercal}$ is the spinor for the annihilation operators. The parameter $V_{j}^{\mathrm{M}}$ controls the type-I and type-II mosaic lattice, which respectively given by
@@
Appendix D 1D spinless QP model
In this section, we show that the 1D spinless QP model with nearest neighbour hopping can also be unified within the spinful QP systems, under the Majorana representation. For a generic 1D spinless QP chain model, the tight binding Hamiltonian is given by
The $t_{j}$ and $V_{j}$ are the hopping coefficients and on-site potential, respectively. Such 1D spinless QP chain can be recasted into the spinful QP chain by introducing the Majorana representation
$c_{j}=\frac{1}{2}(\gamma_{B,j}+i\gamma_{A,j}),\quad\quad c_{j}^{\dagger}=\frac{1}{2}(\gamma_{B,j}-i\gamma_{A,j}),$ (88)
with $\gamma_{\sigma,j}=\gamma_{\sigma,j}^{\dagger}$ and $\{\gamma_{\sigma,j},\gamma_{\sigma;,j^{\prime}}\}=2\delta_{\sigma\sigma^{\prime}}\delta_{jj^{\prime}}$. In this basis, the Hamiltonian becomes
$+$ $\sum_{j}\frac{V_{j}}{2}(1+i\gamma_{B,j}\gamma_{A,j}).$
$\Pi_{j}=\frac{t_{j}}{4}\sigma_{y},\quad M_{j}=\frac{V_{j}}{4}\sigma_{y}.$ (90)
One can notice that, its characteristic is that both hopping coupling and on-site matrix shares the same Pauli component $\sigma_{y}$. This formalism unifies all the well known 1D spinless QP model in this context, including the AA model Aubry and André (1980), the extended AA (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015), the generalized AA (GAA) model Ganeshan et al. (2015), or the GPD model in some context. These are summarized in the table. 3.
Parameters Models
$t_{j}=t\quad V_{j}=2VV_{j}^{\mathrm{d}}$ AA model
$t_{j}=(t+\mu)V_{j}^{\mathrm{od}}\quad V_{j}=2V_{0}V_{j}^{\mathrm{d}}$ EAA model
$t_{j}=t\quad V_{j}=2V_{0}V_{j}^{\mathrm{d}}/(1-aV_{j}^{\mathrm{d}})$ GAA model
Table 3: The parameters of different 1D spinless QP model
@@
In this setup, the ratio $\alpha=k_{s}/k_{p}\approx 0.83$, which is experimentally feasible for observing MEs.
Appendix E Experimental Scheme
@@
where $\phi_{1}$ (without the prime) is the initial phase of the incident beam, and $\phi_{1}^{\prime}$ is an additional phase acquired by ${\bf E}_{1}$ before being reflected back to the atoms. Given that the modulus of the wave vector is $k_{1}=2\pi/\lambda_{1}$, and ignoring the overall constant energy shift, the standing wave field ${\bf E}_{1}$ induces a spin-independent potential given by
${\cal V}_{1}(z)$ $=\frac{V_{1}}{2}\cos(2k_{1}z-\phi_{1}^{\prime}),$
$\text{where}\quad V_{1}$ $=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,10}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (93)
@@
${\cal V}_{p}(z)$ $=\frac{V_{2}}{2}\cos(2k_{2}z-\phi_{2}^{\prime}),$
$\text{where}\quad V_{2}$ $=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,20}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (96)
Then the standing wave ${\bf E}_{3}$ and the traveling wave ${\bf E}_{4}$ together generate a periodic Raman potential along the z-direction:
${\cal V}_{p}(z)$ $=M_{34}e^{{\rm i}(\phi_{4}-\phi_{3}-\phi_{3}^{\prime}/2)}\cos(k_{3}z-\phi_{3}^{\prime}/2),$
$\text{where}\quad M_{34}$ $=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\Omega_{F^{\prime},m_{F}^{\prime};\uparrow,30}^{(J)*}\Omega_{F^{\prime},m_{F}^{\prime};\downarrow,4-}^{(J)}}{\sqrt{2}\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (97)
Here, we adopt the following gauge convention for the spherical basis: $\hat{e}_{z}=-1(\hat{e}_{+}-\hat{e}_{-})/\sqrt{2}$ and $\hat{e}_{y}={\rm i}(\hat{e}_{+}+\hat{e}_{-})/\sqrt{2}$. For simplicity, we choose ‘tune-out’ wavelength $\lambda_{3}\approx 790\,{\rm nm}$ Wen et al. (2021). As a result, the lattice depth generated by ${\bf E}_{3}$ is nealy zero, since the optical shifts from the $D_{1}$ and $D_{2}$ lines cancel each other out footnote_Exp . By comparing this expression with Eq. (E), the following identities hold:
$k_{3}=2k_{2},\quad M_{34}=V_{2}/2,\quad\phi_{3}^{\prime}/2=\phi_{2}^{\prime}.$ (98)

### v4

@@ Appendix E Exact solvability of QP systems from local constraint
Now we show the exact solvability for the spinful quasiperiodic system when the hopping coupling matrices $\Pi_{j}$ or its dual counterpart $\Pi_{n}$ are degenerate
$\det|\Pi_{j}|=0\quad{\rm or}\quad\det|\Pi_{n}|=0,$ (91)
@@
which describe the effective motion with nearest-neighbor hopping and the effective on-site potential is energy dependently dressed by the local modulation from the accompanied internal degree of freedom.
@@
Appendix F 1D QP mosaic model
In this section, we introduce the one-dimensional quasiperiodic mosaic model Wang et al. (2020a); Zhou et al. (2023) and rewrite it into our 1D spinful QP framework. The generic QP mosaic lattice refers that the modulation of on-site potential and hopping coefficients are in a mosaic manner, namely the the QP modulation or the uniform modulation only appears every $\kappa$ sites, as we will see in the following.
The original type-I and type-II QP mosaic model is distinguished by the presence of alternating mosaic pattern in hopping coefficients. If the QP modulation only exhibits in the on-site potential, then the mosaic models belongs to type-I, with Hamiltonian  Wang et al. (2020a)
@@
where the on-site potential is given by
$V_{j}=\begin{cases}V_{0}\cos(2\pi\alpha j+\theta),&j=0\quad{\rm mod}\leavevmode\nobreak\ \kappa,\\ 0,&j\neq 0\quad{\rm mod}\leavevmode\nobreak\ \kappa.\end{cases}$ (102)
Here $c_{j}$ is the annihilation operator at site $j$, and $\lambda$, $V$, and $\theta$ denote the nearest-neighbor hopping coefficient, on-site potential amplitude, and phase offset, respectively. $\alpha$ is an irrational number, and $\kappa$ is an integer. The QP potential periodically occurs every $\kappa$ sites.
If the system exhibits QP mosaic modulation in hopping coefficients, then it belongs to type-II QP mosaic model, with Hamiltonian Zhou et al. (2023)
@@
The conditions for the different states of the two QP mosaic models are summarized in Table 5, and the MEs separating different types of states can be obtained accordingly. The critical state is absent in type-I model and the extended state is absent in type-II model. The critical states in the Type II QP mosaic model are generated by the IDZs in the hopping coefficients Zhou et al. (2023).
Extended Critical Localized
Type I $E<|\lambda^{2}/V_{0}|$ none $E>|\lambda^{2}/V_{0}|$
Type II none $E<|\lambda|$ $E>|\lambda|$
Table 5: The criteria for different states of type I and type II mosaic models. Here $E$ is the eigenenergies of the system, $\lambda$ and $V_{0}$ are the coefficients defined in the main text. The critical state is absent in type I model and the extended state is absent in type II model.
If we focus on the case $\kappa=2$, then both two QP mosaic lattice models exhibit a even-odd sublattice structure, which can be rewritten as the spinful QP chain model. We can relabel the even and odd site as spin-up and spin-down degrees of freedom and further relabel the site index, then the Hamiltonian can be rewritten as
@@
where $c_{j}=(c_{j,\uparrow},c_{j,\downarrow})^{\intercal}$ is the spinor for the annihilation operators. The parameter $V_{j}^{\mathrm{M}}$ controls the type-I and type-II QP mosaic lattice, which respectively given by
@@
Appendix G Experimental Scheme
@@
where $\phi_{1}$ is the initial phase of the incident beam, and $\phi_{1}^{\prime}$ is an additional phase acquired by ${\bf E}_{1}$ before being reflected back to the atoms. Given that the modulus of the wave vector is $k_{1}=2\pi/\lambda_{1}$, and ignoring the overall constant energy shift, the standing wave field ${\bf E}_{1}$ induces a spin-independent potential given by
${\cal V}_{1}(z)=\frac{V_{1}}{2}\cos(2k_{1}z-\phi_{1}^{\prime}),$ (111)
where
$V_{1}=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,10}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (112)
@@
${\cal V}_{p}(z)=\frac{V_{2}}{2}\cos(2k_{2}z-\phi_{2}^{\prime}),$ (115)
where
$V_{2}=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\absolutevalue{\Omega_{F^{\prime},m_{F}^{\prime};\sigma,20}^{(J)}}^{2}}{\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (116)
Then the standing wave ${\bf E}_{3}$ and the traveling wave ${\bf E}_{4}$ together generate a periodic Raman potential along the z-direction
${\cal V}_{p}(z)=M_{34}e^{{\rm i}(\phi_{4}-\phi_{3}-\phi_{3}^{\prime}/2)}\cos(k_{3}z-\phi_{3}^{\prime}/2),$ (117)
with $M_{34}$ given by
$M_{34}=\sum_{J,F^{\prime},m_{F}^{\prime}}\frac{\Omega_{F^{\prime},m_{F}^{\prime};\uparrow,30}^{(J)*}\Omega_{F^{\prime},m_{F}^{\prime};\downarrow,4-}^{(J)}}{\sqrt{2}\Delta_{F^{\prime},m_{F}^{\prime}}^{(J)}}.$ (118)
Here, we adopt the following gauge convention for the spherical basis: $\hat{e}_{z}=-1(\hat{e}_{+}-\hat{e}_{-})/\sqrt{2}$ and $\hat{e}_{y}={\rm i}(\hat{e}_{+}+\hat{e}_{-})/\sqrt{2}$. For simplicity, we choose ‘tune-out’ wavelength $\lambda_{3}\approx 790\,{\rm nm}$ Wen et al. (2021). As a result, the lattice depth generated by ${\bf E}_{3}$ is nearly zero, since the optical shifts from the $D_{1}$ and $D_{2}$ lines cancel each other out footnote_Exp . By comparing this expression with Eq. (115), the following identities hold:
$k_{3}=2k_{2},\ M_{34}=V_{2}/2,\ \phi_{3}^{\prime}/2=\phi_{2}^{\prime}.$ (119)